# Foundation Model of Electronic Medical Records for Adaptive Risk Estimation

### v3

@@ 3 Results
Following the tokenization process, the data of 299,721 unique patients from the MIMIC-IV dataset was converted into 285,622 PHTs, which were subsequently used for training and testing. The discrepancy arises from some patients lacking associated data after tokenization. Of the total PHTs, approximately 60% (180,733) contained hospital admissions records. The tokenized dataset comprised over 357 million tokens in total. Detailed information regarding the MIMIC-IV data used, patient demographics, characteristics of the PHTs and tokens, and descriptive statistics are provided in supplementary data (Table S2,S7,S6,LABEL:tab:detailed-token-stats). The model was trained and validated on 90% of the PHTs, with the remaining 10% reserved for testing. During inference, at least 100 fPHTs were generated for each investigated task.
The predictive performance of ETHOS and MEDS-Tab was evaluated for four critical clinical outcomes: hospital mortality, ICU admission, prolonged hospital stay, and a composite risk score (HM+IA+PS). Prolonged stay was defined as a stay longer than 90th percentile of all stays. All predictions were performed at patient admission. As summarized in Figure 4, 5, and Table S1, ETHOS consistently outperformed MEDS-Tab across all outcome measures, demonstrating superior AUC values. Notably, ETHOS yielded higher AUC values across all racial groups, with the most significant improvement observed among Asian and Hispanic patients. The model’s robustness across diverse populations suggests its potential for mitigating disparities in predictive accuracy.
@@
In addition to risk which are part of ARES and to contextualize the predictive capabilities of ETHOS, we compared its performance against traditional early warning scores and other ML models. Figure 5 presents the AUC values (ROC curves in Figure S2) for key ED benchmark tasks: hospitalization at triage, critical outcomes within 12 hours of triage, and ED re-presentation within 72 hours post-discharge. ETHOS demonstrated consistently superior predictive accuracy across all evaluated tasks. We provide detailed numerical values in supplementary data (Table S3,S4,S5).
@@ 4 Discussion
The ARES framework introduces an innovative approach to building predictive models by leveraging cutting-edge artificial intelligence technology. Several aspects of this approach distinguish it from traditional models. First, ARES enables dynamic risk estimation at any time during a patient’s stay, from admission to discharge. Powered by ETHOS [Renc2024-jf], ARES utilizes PHTs and incorporates all available clinical information at the time of risk estimation. Unlike traditional models, which rely on static data points such as information collected within 24 hours after admission or ED presentation or data up to triage [Xie2022-ur, Meng2022-bk], ARES continuously adapts to the patient’s evolving clinical status. This adaptability overcomes a key limitation of static models, which may not perform optimally outside the narrow time frames for which they are designed. This capability is demonstrated in the accompanying Figure 3 and LABEL:tab:timeline-examples, which illustrate how risk evolves over time during a patient’s hospital stay. These visualizations, which depict how personalized risk evolves over time to reach the current estimates, provide insights into the specific factors driving model predictions for each patient. They highlight clinical events associated with increased or decreased risk, offering real-time explainability. By identifying the most influential features contributing to an individual’s risk assessment, ARES empowers clinicians with a clearer understanding of the rationale behind each prediction.
@@
As illustrated in Figure 1, ARES can estimate risk for various critical events, such as in-hospital mortality, ICU admission, and prolonged hospital stays. Beyond these standard metrics, additional indicators can be integrated seamlessly, including the risk of ICU admission during a specific length of stay, ICU readmission, acute kidney injury, sepsis, cardiac arrest, or 30-day readmission, and others. The ETHOS model, which underpins ARES, allows for the dynamic combination of these risks into composite measures while accounting for their interdependencies. For example, the occurrence of mortality on Day 8 would render the probability of a 10-day hospital stay zero. This ability to incorporate conditional and causal relationships between tracked events is another strength of ARES. Importantly, integrating additional metrics does not require model retraining or modifications to the ETHOS model. Once a range of possible future PHTs has been generated, any additional metrics can be calculated with minimal computational resources, making ARES scalable and adaptable to diverse healthcare settings.
@@
In its current implementation, ETHOS distills multiple fPHTs into a single predictive decision, such as inpatient mortality. However, this approach overlooks the wealth of longitudinal information contained in these trajectories, including the sequence of clinical events that lead to a particular outcome, or the absence thereof. By merely predicting the likelihood of an adverse event, valuable insights into the pathways that contribute to deterioration or recovery remain underutilized. Expanding ETHOS to provide a more granular, trajectory-based interpretation of risk would allow clinicians not only to assess a patient’s probability of experiencing a critical event but also to understand the evolving clinical course leading to that outcome including the cost. This enhanced approach would address a key limitation highlighted in the early warning paradox [Logan-Ellis2025-qr], where models trained on retrospective data may fail to capture the full complexity of clinical interventions and their effects on patient outcomes. Moving forward, we aim to refine ETHOS to incorporate and visualize these probabilistic trajectories. This will equip clinicians with deeper, more actionable insights into clinical risk dynamics and potentially provide new information about causality in patient outcomes.
This study has limitations. ETHOS was demonstrated using PHTs derived from MIMIC-IV-ED data, and its direct applicability to data from other institutions may be limited without retraining using additional data from other institutions. Electronic medical record (EMR) systems vary significantly across institutions, influenced by differences in clinical practice, care pathways, patient populations, and geographic location. These variations can impede the direct transferability of AI models trained on one dataset to another. In certain applications, such as radiology or pathology, data inputs like medical images are relatively standardized, allowing models trained in one institution to perform well in others. However, EMR data pose unique challenges due to their variability. Models trained on data from one institution may produce inaccurate risk estimates when applied to data from another, particularly if clinical practices differ [Rajkomar2019-jk]. To mitigate this limitation, the model code for ETHOS-ARES is compatible with the MEDS [Arnrich2024-rc] health AI data standard, making it easier for other researchers to reliably train identical model architectures on their local data.
@@
Data standardization is often proposed as a solution to address the challenges of variability in healthcare data. However, achieving meaningful standardization would require identifying commonalities between healthcare systems, an endeavor that may not be feasible given the diversity of clinical practices, patient populations, and institutional workflows. In our view, a more robust approach is to train AI models, such as ETHOS, on raw data from diverse institutions, allowing the model itself to learn and interpret the underlying patterns and clinical pathways. This approach mirrors the capability of large language models (LLMs) to discern meaning from vastly different styles of text and presentations or even different languages, leveraging the same foundational transformer architecture as ETHOS.
@@
In summary, artificial intelligence advancements have unlocked unprecedented opportunities for innovative solutions like ARES, which leverage large amounts of heterogeneous data to develop general-purpose models with superior predictive power compared to state-of-the-art methods. ARES not only enables dynamic, personalized risk estimation but also provides real-time explainability, empowering clinicians to make more informed decisions. Furthermore, its modular design and the underlying ETHOS model allow for seamless integration of additional data types, such as radiology, genetics, and other institutional datasets, paving the way for even greater predictive accuracy and applicability in diverse healthcare settings.
@@ Acknowledgments
We thank Kinga Renc, M.Arch, for her invaluable assistance with graphic design. This work was supported in part by National Institutes of Health (NIH) grant number HL159183.
[…]
@@
Table S1: ETHOS performance on the ARES tasks with a breakdown for demographic subgroups. This table presents the predictive performance (AUROC with 95% confidence intervals) of ETHOS (top) and MEDS-Tab (bottom) for four critical clinical outcomes used in ARES: Hospital Mortality, ICU Admission, Prolonged Hospital Stay (>10 days), and a Composite Risk Score (HM+IA+PS). The prevalence rates of each outcome are provided for reference. Performance metrics are further stratified by gender and race to assess potential disparities in model performance across demographic subgroups.
@@
      Asian 0.953 [0.947, 0.965] 0.939 [0.938, 0.955] 0.871 [0.854, 0.878] 0.913 [0.907, 0.926]
      Black 0.959 [0.954, 0.970] 0.927 [0.926, 0.939] 0.867 [0.860, 0.877] 0.899 [0.888, 0.910]
      Hispanic 0.942 [0.933, 0.981] 0.940 [0.934, 0.948] 0.872 [0.849, 0.883] 0.909 [0.893, 0.919]
      Other 0.977 [0.962, 0.990] 0.961 [0.957, 0.969] 0.863 [0.837, 0.883] 0.927 [0.918, 0.940]
      Unknown 0.890 [0.875, 0.910] 0.933 [0.921, 0.941] 0.796 [0.773, 0.799] 0.933 [0.927, 0.947]
      White 0.920 [0.909, 0.925] 0.918 [0.915, 0.920] 0.840 [0.838, 0.843] 0.890 [0.888, 0.893]
@@
      Asian 0.895 [0.825, 0.909] 0.916 [0.906, 0.920] 0.819 [0.795, 0.847] 0.877 [0.860, 0.885]
      Black 0.918 [0.892, 0.948] 0.920 [0.909, 0.928] 0.836 [0.824, 0.836] 0.874 [0.862, 0.883]
      Hispanic 0.890 [0.846, 0.954] 0.922 [0.901, 0.927] 0.851 [0.839, 0.853] 0.900 [0.880, 0.907]
      Other 0.933 [0.889, 0.945] 0.943 [0.938, 0.965] 0.830 [0.804, 0.854] 0.915 [0.902, 0.930]
      Unknown 0.789 [0.779, 0.824] 0.953 [0.950, 0.958] 0.750 [0.705, 0.783] 0.926 [0.922, 0.942]
      White 0.871 [0.868, 0.878] 0.907 [0.904, 0.909] 0.806 [0.798, 0.807] 0.867 [0.865, 0.867]
Table S2: Demographic characteristics of the dataset analyzed in this study. Performance Comparison of ETHOS and MEDS-Tab Across Clinical Outcomes and Demographic Subgroups. This table presents the predictive performance (AUROC with 95% confidence intervals) of ETHOS (top) and MEDS-Tab (bottom) for four critical clinical outcomes used in ARES: Hospital Mortality, ICU Admission, Prolonged Hospital Stay (>10 days), and a Composite Risk Score (HM+IA+PS). The prevalence rates of each outcome are provided for reference. Performance metrics are further stratified by gender and race to assess potential disparities in model performance across demographic subgroups.
@@
Table S3: Prediction of Hospitalization At Triage. Performance comparison of various models for predicting hospitalization at triage, evaluated using AUROC, AUPRC, sensitivity, and specificity (95% confidence intervals in brackets). The thresholds for sensitivity and specificity were determined by finding the operating point on the ROC curve closest to (0,1). ETHOS demonstrates superior performance across all metrics, achieving the highest AUROC (0.912), AUPRC (0.887), sensitivity (0.849), and specificity (0.820), outperforming all other methods, including traditional scoring systems and machine learning models.
@@
LR 0.809 [0.807, 0.813] 0.773 [0.770, 0.780] 0.751 [0.729, 0.757] 0.723 [0.720, 0.747]
RF 0.817 [0.814, 0.821] 0.785 [0.778, 0.791] 0.767 [0.735, 0.773] 0.716 [0.714, 0.748]
GB 0.819 [0.816, 0.822] 0.792 [0.787, 0.797] 0.753 [0.732, 0.771] 0.728 [0.712, 0.750]
MLP 0.823 [0.821, 0.827] 0.797 [0.792, 0.803] 0.748 [0.740, 0.778] 0.740 [0.720, 0.752]
ESI 0.712 [0.708, 0.716] 0.632 [0.628, 0.639] 0.584 [0.577, 0.591] 0.784 [0.782, 0.790]
NEWS 0.581 [0.577, 0.585] 0.555 [0.548, 0.559] 0.563 [0.554, 0.568] 0.546 [0.542, 0.552]
NEWS2 0.565 [0.561, 0.569] 0.538 [0.532, 0.543] 0.519 [0.510, 0.526] 0.570 [0.567, 0.576]
REMS 0.666 [0.661, 0.669] 0.605 [0.599, 0.610] 0.605 [0.553, 0.717] 0.641 [0.544, 0.712]
MEWS 0.558 [0.555, 0.561] 0.521 [0.516, 0.527] 0.296 [0.292, 0.299] 0.812 [0.807, 0.817]
CART 0.673 [0.668, 0.676] 0.617 [0.609, 0.622] 0.703 [0.699, 0.707] 0.578 [0.571, 0.583]
Med2Vec 0.815 [0.812, 0.818] 0.779 [0.774, 0.783] 0.741 [0.732, 0.756] 0.739 [0.726, 0.754]
AutoScore 0.794 [0.791, 0.797] 0.755 [0.750, 0.761] 0.745 [0.714, 0.749] 0.698 [0.692, 0.730]
MEDS-Tab 0.846 [0.842, 0.850] 0.862 [0.859, 0.866] 0.723 [0.713, 0.729] 0.807 [0.801, 0.819]
ETHOS (ours) 0.912 [0.909, 0.915] 0.887 [0.881, 0.890] 0.849 [0.846, 0.852] 0.820 [0.817, 0.823]
Table S4: Prediction of Critical Outcome Within 12h At Triage. Performance comparison of various models for predicting critical outcomes within 12 hours of triage, evaluated using AUROC, AUPRC, sensitivity, and specificity (95% confidence intervals in brackets). The thresholds for sensitivity and specificity were determined by finding the operating point on the ROC curve closest to (0,1). ETHOS achieves the highest performance across most of the metrics, with an AUROC of 0.937, AUPRC of 0.649, sensitivity of 0.858, and specificity of 0.863, substantially outperforming all other methods, including traditional scoring systems and machine learning models.
@@
LR 0.869 [0.864, 0.872] 0.327 [0.309, 0.338] 0.804 [0.795, 0.822] 0.777 [0.763, 0.784]
RF 0.875 [0.868, 0.881] 0.385 [0.368, 0.405] 0.813 [0.789, 0.825] 0.785 [0.783, 0.809]
GB 0.884 [0.880, 0.887] 0.404 [0.386, 0.415] 0.804 [0.799, 0.829] 0.799 [0.774, 0.800]
MLP 0.888 [0.883, 0.891] 0.397 [0.378, 0.407] 0.817 [0.803, 0.838] 0.795 [0.782, 0.812]
ESI 0.807 [0.802, 0.812] 0.198 [0.188, 0.207] 0.877 [0.870, 0.888] 0.640 [0.637, 0.644]
NEWS 0.638 [0.625, 0.654] 0.155 [0.144, 0.165] 0.458 [0.439, 0.480] 0.798 [0.796, 0.803]
NEWS2 0.621 [0.609, 0.635] 0.141 [0.132, 0.149] 0.596 [0.404, 0.612] 0.536 [0.535, 0.826]
REMS 0.679 [0.668, 0.689] 0.109 [0.102, 0.117] 0.674 [0.656, 0.693] 0.605 [0.602, 0.608]
MEWS 0.621 [0.610, 0.631] 0.112 [0.106, 0.117] 0.442 [0.421, 0.461] 0.774 [0.770, 0.778]
CART 0.706 [0.695, 0.713] 0.154 [0.142, 0.161] 0.591 [0.569, 0.603] 0.723 [0.719, 0.727]
Med2Vec 0.874 [0.870, 0.878] 0.334 [0.313, 0.347] 0.825 [0.794, 0.838] 0.767 [0.761, 0.797]
AutoScore 0.857 [0.853, 0.861] 0.311 [0.294, 0.319] 0.789 [0.753, 0.813] 0.762 [0.745, 0.794]
MEDS-Tab 0.814 [0.805, 0.822] 0.420 [0.403, 0.440] 0.776 [0.761, 0.790] 0.663 [0.658, 0.667]
ETHOS (ours) 0.937 [0.932, 0.943] 0.649 [0.628, 0.666] 0.858 [0.850, 0.867] 0.863 [0.856, 0.869]
Table S5: Prediction of Emergency Department Re-presentation Within 72h. Performance comparison of various models for predicting emergency department re-presentation within 72 hours, evaluated using AUROC, AUPRC, sensitivity, and specificity (95% confidence intervals in brackets). The thresholds for sensitivity and specificity were determined by finding the operating point on the ROC curve closest to (0,1). ETHOS demonstrates superior performance, achieving the highest AUROC (0.740), AUPRC (0.199), sensitivity (0.659), and specificity (0.696), outperforming all other methods and showcasing its effectiveness for this challenging task.

### v4

@@ 3 Results
Following the tokenization process, the data of 299,721 unique patients from the MIMIC-IV dataset was converted into 285,622 PHTs, which were subsequently used for training and testing. Patients were excluded if they had no usable data after tokenization. This occurs when patients in MIMIC have little or no structured information available, or when the available information (e.g., clinical notes, imaging) is not tokenized in the current ETHOS version. Of the total PHTs, approximately 63% (180,733) contained hospital admissions records. The tokenized dataset comprised over 360 million tokens in total. In the Supplementary Materials, we provide detailed information regarding the MIMIC-IV data used (Table S7), patient demographics (Table S2), characteristics of the PHTs (Table S6) and tokens (LABEL:tab:detailed-token-stats). The model was trained and validated on 90% of the PHTs, with the remaining 10% reserved for testing. During inference, at least $N=100$ fPHTs were generated for each investigated task.
The predictive performance of ARES and MEDS-Tab was evaluated on three individual clinical endpoints—hospital mortality (HM), ICU admission (IA), and prolonged hospital stay (PS; defined as length of stay >90th percentile)—and, as an illustrative demonstration of joint risk modeling, on a composite criterion combining these events (HM-IA-PS). The prevalence of these tasks is: 1.85%, 15.44%, 9.01% and 20.39%, respectively. This composite endpoint demonstrates ARES’s capacity to compute joint probabilities across heterogeneous outcomes and to naturally model their statistical dependencies. The composite score represents the cumulative risk of clinician-defined critical events. All predictions were generated at the hospital admission. As summarized in Figure 5, Figure 4, and Table S1, ARES consistently outperformed MEDS-Tab across both individual and composite endpoints, achieving higher AUC values in every case. Notably, these gains were observed across all racial subgroups, with the most pronounced improvements for Asian and Hispanic patients, indicating ARES’s robustness and its potential to reduce disparities in predictive accuracy.
@@
In addition to risk which are part of ARES and to contextualize the predictive capabilities of ETHOS, we compared its performance against traditional early warning scores and other ML models. Figure 5 presents the AUC values (ROC curves in Figure S2) for key ED benchmark tasks: hospitalization at triage, critical outcomes within 12 hours of triage, and ED re-presentation within 72 hours post-discharge. ETHOS demonstrated consistently superior predictive accuracy across all evaluated tasks. We provide detailed numerical values in the Supplement (Table S3,S4,S5).
@@ 4 Discussion
The ARES framework introduces an innovative approach to building predictive models by leveraging cutting-edge artificial intelligence technology. Several aspects of this approach distinguish it from traditional models. First, ARES enables dynamic risk estimation at any time during a patient’s stay, from admission to discharge. Powered by ETHOS [15], ARES utilizes PHTs and incorporates all available clinical information at the time of risk estimation. Unlike traditional models, which rely on static data points such as information collected within 24 hours after admission or ED presentation or data up to triage [29, 41], ARES continuously adapts to the patient’s evolving clinical status. This adaptability overcomes a key limitation of static models, which may not perform optimally outside the narrow time frames for which they are designed. This capability is demonstrated in the accompanying Figure 3 and LABEL:tab:timeline-examples, which illustrate how risk evolves over time during a patient’s hospital stay. These visualizations, which depict how personalized risk evolves over time to reach the current estimates, provide insights into the specific factors driving model predictions for each patient. They highlight clinical events associated with increased or decreased risk, offering real-time explainability. By identifying the most influential features contributing to an individual’s risk assessment, ARES have the potential to empower clinicians with a clearer understanding of the rationale behind each prediction.
@@
As illustrated in Figure 1, ARES can estimate risk for various critical events, such as in-hospital mortality, ICU admission, and prolonged hospital stays. Beyond these standard metrics, additional indicators can be integrated seamlessly, including the risk of ICU admission during a specific length of stay, ICU readmission, acute kidney injury, sepsis, cardiac arrest, or 30-day readmission, and others. The ETHOS model, which underpins ARES, allows for the dynamic combination of these risks into composite measures while accounting for their interdependencies. For example, the occurrence of mortality on day 8 would render the probability of a 10-day hospital stay zero. This ability to incorporate conditional and causal relationships between tracked events is another strength of ARES. Importantly, integrating additional metrics does not require model retraining or modifications of ETHOS. Once a range of possible future PHTs has been generated, any additional metrics can be calculated with minimal computational resources, making ARES scalable and adaptable to diverse healthcare settings.
@@
In its current implementation, ETHOS distills multiple fPHTs into a single predictive decision, such as inpatient mortality. However, this approach overlooks the wealth of longitudinal information contained in these trajectories, including the sequence of clinical events that lead to a particular outcome, or the absence thereof. By merely predicting the likelihood of an adverse event, valuable insights into the pathways that contribute to deterioration or recovery remain underutilized. Expanding ARES to provide a more granular, trajectory-based interpretation of risk would allow clinicians not only to assess a patient’s probability of experiencing a critical event but also to understand the evolving clinical course leading to that outcome including the cost. This enhanced approach would address a key limitation highlighted in the early warning paradox [42], where models trained on retrospective data may fail to capture the full complexity of clinical interventions and their effects on patient outcomes. Moving forward, we aim to refine ARES to incorporate and visualize these probabilistic trajectories. This will equip clinicians with deeper, more actionable insights into clinical risk dynamics and potentially provide new information about causality in patient outcomes.
We recognize that, to date, ARES has not undergone formal usability testing with frontline clinicians, yet their ultimate impact depends on seamless integration into real-world workflows. Emergency medicine specialists on our team have provided informal feedback on feasibility and clarity of the risk estimates and explanatory highlights, and we are now designing pilot simulations in which physicians will “round” on de-identified patient cases presented through mock electronic charts powered by ARES. These studies, first leveraging MIMIC-derived timelines and subsequently our own Mass General Brigham data, will allow us to observe decision points, gather qualitative feedback on timing and interpretability of alerts, and refine both the user interface and explanation formats. We anticipate that iterative, case-based testing will guide the development of a clinician-centered dashboard, ensuring that ARES’s predictions align with care priorities and support timely, actionable insights in the emergency setting.
This study has several important limitations. First, although we demonstrated ETHOS using PHTs derived from the MIMIC-IV-ED dataset, its performance on data from other institutions may be compromised without retraining on external cohorts. Electronic health record systems and clinical workflows differ substantially across hospitals—driven by variations in documentation practices, patient case-mix, and care protocols—so models trained on one site can yield misleading risk estimates when deployed elsewhere. Moreover, our training data may harbor demographic and institutional biases (for example, overrepresentation of certain age, race, or socioeconomic groups), which could impair generalizability and exacerbate health inequities if unaddressed. The MIMIC dataset is relatively small, and although it contains dense, diverse information, the limited cohort size inevitably constrains generalization. Our model is explicitly designed to scale, and we expect its performance to further improve when trained on larger and more diverse datasets that capture a broader range of clinical variability. We have not yet conducted a thorough fairness audit to quantify potential disparities in ETHOS’s predictions across sex, race, or ethnicity. By contrast, in domains such as radiology or pathology, data inputs like images are relatively standardized, enabling easier cross-institution transfer. To facilitate broader validation and retraining, we have ensured that the ETHOS-ARES codebase is fully compatible with the MEDS health AI data standard [28]. This interoperability simplifies the process for other researchers to apply identical model architectures to their local data, perform bias and subgroup analyses, and iteratively refine ETHOS for diverse patient populations.
We also recognize that our evaluation on the extensively curated MIMIC-IV dataset may underestimate challenges encountered in real-world EHRs, which often exhibit higher rates of missing or irregular data, temporal shifts in documentation and care processes, and evolving patient populations. Although ETHOS is designed to operate on incomplete timelines, elevated missingness will still impair model accuracy. It is unclear if temporal biases or practice changes may introduce drift over time for ETHOS because the inferences are based on the context which contemporary to predictions, but this has not been investigated yet. Future work will systematically assess ETHOS’s resilience to these factors and develop strategies for ongoing recalibration in heterogeneous clinical environments. In addition, our current benchmarking employed only classical risk-prediction methods. While sufficient to demonstrate ETHOS’s competitiveness on established tasks, more sophisticated approaches could yield higher benchmark scores. Even if such models were to match or slightly exceed our performance on static benchmarks, this would not diminish ETHOS’s primary contribution: a dynamic, explainable framework that adapts predictions in real time as new clinical data become available. Future work will expand benchmarking to include a broader set of advanced methods (e.g. [22]).
In current implementation, we exclude unstructured clinical text and because of that ETHOS may miss nuanced patient information, such as narrative impressions or social determinants, that could enhance risk estimation and zero-shot generalizability. Integrating free-text notes poses challenges in segmenting and embedding variable-length narratives alongside structured events without overwhelming the model’s capacity. In future work, we will explore the use of pretrained clinical-language-model embeddings, hierarchical chunking of note content, and multimodal fusion techniques to incorporate these rich data into the PHTs.
@@
Data standardization is often proposed as a solution to address the challenges of variability in healthcare data. However, achieving meaningful standardization would require identifying commonalities between healthcare systems, an endeavor that may not be feasible given the diversity of clinical practices, patient populations, and institutional workflows. An alternative is to train AI models, such as ETHOS, on raw data from diverse institutions, allowing the model itself to learn and interpret the underlying patterns and clinical pathways. This approach mirrors the capability of large language models to discern meaning from vastly different styles of text and presentations or even different languages, leveraging the same transformer architecture as ETHOS. We performed an energy consumption analysis comparing training of ETHOS to other known LLMs that can be seen in Table S9.
@@
In summary, recent advances in AI have created unprecedented opportunities for innovative solutions like ARES, which harness large volumes of heterogeneous data to build general-purpose models whose predictive performance exceeds state-of-the-art methods. ARES delivers dynamic, personalized risk estimates and offers real-time explainability, empowering clinicians to make better-informed decisions. Moreover, its modular architecture and the underlying ETHOS framework enable seamless integration of additional data modalities, such as radiology, genomics, and other institutional datasets, further enhancing predictive accuracy and broadening applicability across diverse healthcare environments. Although our results are promising, the clinical impact remains uncertain. Demonstrating ARES’s true utility in real-world settings will be the focus of our future work.
@@ Acknowledgments
We thank Kinga Renc, M.Arch, for her invaluable assistance with graphic design. We also thank Ethan Steinberg for the careful and very helpful review of our codebase. This work was supported in part by National Institutes of Health (NIH) grant number HL159183.
@@
Table S1: ETHOS performance on ARES tasks with a breakdown for demographic subgroups. This table presents the predictive performance (AUROC with 95% confidence intervals) of ETHOS (top) and MEDS-Tab (bottom) for four critical clinical outcomes used in ARES: Hospital Mortality, ICU Admission, Prolonged Hospital Stay (>10 days), and a Composite Risk Score (HM+IA+PS). The prevalence rates of each outcome are provided for reference. Performance metrics are further stratified by gender and race to assess potential disparities in model performance across demographic subgroups.
@@
    Asian 0.954 [0.925, 0.978] 0.944 [0.930, 0.958] 0.880 [0.858, 0.900] 0.921 [0.905, 0.936]
    Black 0.955 [0.935, 0.972] 0.933 [0.924, 0.943] 0.872 [0.858, 0.885] 0.906 [0.897, 0.915]
    Hispanic 0.964 [0.929, 0.987] 0.936 [0.918, 0.951] 0.881 [0.856, 0.903] 0.910 [0.893, 0.927]
    Other 0.985 [0.975, 0.993] 0.956 [0.937, 0.972] 0.862 [0.831, 0.889] 0.932 [0.915, 0.949]
    Unknown 0.886 [0.852, 0.911] 0.941 [0.928, 0.953] 0.790 [0.757, 0.819] 0.942 [0.928, 0.954]
    White 0.928 [0.918, 0.937] 0.923 [0.919, 0.927] 0.844 [0.838, 0.851] 0.897 [0.892, 0.901]
@@
    Asian 0.895 [0.844, 0.938] 0.916 [0.898, 0.933] 0.819 [0.786, 0.849] 0.877 [0.858, 0.897]
    Black 0.918 [0.896, 0.938] 0.920 [0.909, 0.930] 0.836 [0.823, 0.851] 0.874 [0.864, 0.885]
    Hispanic 0.890 [0.795, 0.954] 0.922 [0.904, 0.939] 0.851 [0.826, 0.874] 0.900 [0.882, 0.914]
    Other 0.933 [0.887, 0.964] 0.943 [0.923, 0.959] 0.830 [0.800, 0.860] 0.915 [0.896, 0.933]
    Unknown 0.789 [0.748, 0.829] 0.953 [0.942, 0.964] 0.750 [0.714, 0.780] 0.926 [0.910, 0.940]
    White 0.871 [0.857, 0.885] 0.907 [0.903, 0.912] 0.806 [0.799, 0.814] 0.867 [0.862, 0.872]
Table S2: Demographic characteristics of the dataset analyzed in this study. The table summarizes key demographic attributes of the dataset, stratified into Train/Validation, Test, and Total splits. Patient numbers, mean age (with standard deviation), and distribution across gender, race, and marital status are shown, with percentages provided in parentheses. The data highlights the representation of each subgroup within the splits, providing context for the population characteristics in the dataset.
@@
Table S3: Prediction of Hospitalization At Triage. Performance comparison of various models for predicting hospitalization at triage, evaluated using AUROC, AUPRC, sensitivity, and specificity (95% confidence intervals in brackets). The thresholds for sensitivity and specificity were determined by finding the operating point on the ROC curve closest to (0,1). ETHOS demonstrates superior performance across all metrics, outperforming all other methods, including traditional scoring systems and machine learning models.
@@
LR 0.809 [0.805, 0.813] 0.775 [0.769, 0.781] 0.734 [0.721, 0.750] 0.736 [0.721, 0.749]
Med2Vec 0.816 [0.812, 0.820] 0.782 [0.775, 0.788] 0.751 [0.734, 0.770] 0.728 [0.711, 0.745]
RF 0.817 [0.814, 0.821] 0.785 [0.779, 0.791] 0.759 [0.736, 0.764] 0.726 [0.720, 0.747]
GB 0.819 [0.815, 0.823] 0.792 [0.786, 0.798] 0.753 [0.731, 0.770] 0.728 [0.715, 0.751]
MLP 0.822 [0.818, 0.826] 0.796 [0.790, 0.802] 0.754 [0.743, 0.775] 0.734 [0.716, 0.745]
esi 0.712 [0.707, 0.716] 0.632 [0.625, 0.638] 0.584 [0.577, 0.590] 0.784 [0.779, 0.789]
NEWS 0.581 [0.576, 0.586] 0.555 [0.548, 0.561] 0.563 [0.556, 0.569] 0.546 [0.540, 0.553]
NEWS2 0.565 [0.560, 0.570] 0.538 [0.532, 0.544] 0.519 [0.512, 0.526] 0.570 [0.564, 0.577]
REMS 0.666 [0.661, 0.671] 0.605 [0.598, 0.612] 0.605 [0.552, 0.722] 0.641 [0.545, 0.711]
MEWS 0.558 [0.553, 0.562] 0.521 [0.515, 0.527] 0.296 [0.289, 0.302] 0.812 [0.806, 0.817]
CART 0.673 [0.668, 0.678] 0.617 [0.610, 0.624] 0.703 [0.696, 0.709] 0.578 [0.571, 0.585]
MEDS-Tab 0.863 [0.860, 0.866] 0.879 [0.876, 0.883] 0.746 [0.735, 0.754] 0.820 [0.809, 0.835]
ETHOS (ours) 0.946 [0.944, 0.947] 0.945 [0.943, 0.947] 0.868 [0.859, 0.876] 0.864 [0.856, 0.873]
Table S4: Prediction of Critical Outcome Within 12h At Triage. Performance comparison of various models for predicting critical outcomes within 12 hours of triage, evaluated using AUROC, AUPRC, sensitivity, and specificity (95% confidence intervals in brackets). The thresholds for sensitivity and specificity were determined by finding the operating point on the ROC curve closest to (0,1). ETHOS achieves the highest performance across most of the metrics, substantially outperforming all other methods, including traditional scoring systems and machine learning models.
@@
LR 0.875 [0.868, 0.882] 0.308 [0.288, 0.328] 0.813 [0.792, 0.836] 0.782 [0.766, 0.803]
Med2Vec 0.880 [0.872, 0.887] 0.324 [0.305, 0.346] 0.817 [0.799, 0.852] 0.787 [0.762, 0.804]
RF 0.881 [0.873, 0.888] 0.362 [0.343, 0.386] 0.812 [0.794, 0.829] 0.792 [0.788, 0.796]
GB 0.891 [0.884, 0.897] 0.389 [0.367, 0.412] 0.836 [0.804, 0.848] 0.788 [0.779, 0.812]
MLP 0.892 [0.886, 0.898] 0.372 [0.352, 0.396] 0.845 [0.806, 0.855] 0.784 [0.780, 0.823]
esi 0.821 [0.814, 0.829] 0.190 [0.178, 0.201] 0.900 [0.887, 0.913] 0.637 [0.632, 0.642]
NEWS 0.637 [0.624, 0.651] 0.139 [0.127, 0.154] 0.461 [0.440, 0.483] 0.796 [0.792, 0.801]
NEWS2 0.622 [0.610, 0.636] 0.130 [0.118, 0.144] 0.416 [0.402, 0.605] 0.822 [0.533, 0.826]
REMS 0.672 [0.661, 0.683] 0.093 [0.087, 0.102] 0.662 [0.642, 0.683] 0.602 [0.597, 0.607]
MEWS 0.623 [0.611, 0.634] 0.101 [0.093, 0.110] 0.445 [0.424, 0.466] 0.772 [0.768, 0.776]
CART 0.699 [0.687, 0.710] 0.134 [0.123, 0.147] 0.579 [0.557, 0.600] 0.720 [0.716, 0.725]
MEDS-Tab 0.853 [0.846, 0.861] 0.513 [0.493, 0.531] 0.735 [0.717, 0.752] 0.764 [0.759, 0.771]
ETHOS (ours) 0.945 [0.941, 0.950] 0.696 [0.678, 0.712] 0.876 [0.860, 0.898] 0.873 [0.852, 0.889]
Table S5: Prediction of Emergency Department Re-presentation Within 72h. Performance comparison of various models for predicting emergency department re-presentation within 72 hours, evaluated using AUROC, AUPRC, sensitivity, and specificity (95% confidence intervals in brackets). The thresholds for sensitivity and specificity were determined by finding the operating point on the ROC curve closest to (0,1). ETHOS demonstrates superior performance, outperforming all other methods and showcasing its effectiveness for this challenging task.